# Sustainable Visions: Unsupervised Machine Learning Insights on Global Development Goals

### v1

@@ 3.5 Cluster dynamics
The evolution of the SDGs in each country can be tracked over time using the full 17-dimensional vector of indicators. Each component of this vector is an indicator valued between 0 and 1, representing the extent to which the country is achieving a specific goal. This makes the vector point formed by these 17 values, denoted as $\vec{r}_{i}$, a unique vector that represents the ideal state being pursued. Consequently, for any given country at any point in time, the Euclidean distance to $\vec{r}_{i}$ can be monitored and visualized. Furthermore, by utilizing the clusters described in Table B.1, it is possible to compute the distribution of distances to the ideal point for each year. Figure 8 displays the probability density functions for the six clusters in the years 2000, 2010, and 2020. This figure reveals a continuous shift in the distribution functions, maintaining an approximately Gaussian shape.
The Gaussian like distribution for the distances to the ideal goals within each cluster as seen in Figure  8 suggests that the countries grouped together share a similar pattern in their progress towards the SDGs. Moreover, the fact that these clusters are Gaussian indicates that within each of them, the countries are relatively homogenous in terms of their distance to the ideal SDG foals. This homogeneity implies that the factors influencing their progress are consistent across the countries in each cluster and that naturally group together based on shared challenges or opportunities in achieving the SDGs. These clusters likely reflect countries that, despite their geographical or cultural differences, face similar systemic issues or have adopted similar strategies in their progress toward the SDGs.
@@
Fitted probability density functions for the distribution of distances to the ideal point for three different years and for each cluster on Table B.1. The colors for the clusters are the same as in Figure 5. The Gaussian like distributions observed indicate a central tendency and homogeneity within each cluster with respect to their progress toward the SDGs.
@@
(A) Trajectories of clusters obtained in Table B.1 toward the ideal achievement of the Sustainable Development Goals (SDGs) over time. Each line represents the projected average distance to the ideal point for a cluster of countries, with actual data points marked by dots, tracking progress from 2000 to 2022. (B) Extrapolation of same lines to year 2100. The vertical dotted line at 2030 marks the critical global warming threshold of 1.5°C, a pivotal point that could drastically influence the ability of countries to meet their SDG targets. Although some clusters appear to be on track to achieve the goals several decades later, the potential impacts of crossing this climate threshold, along with other global challenges such as resource depletion, biodiversity loss, and population growth, suggest that these optimistic trajectories may not be realized without significant and urgent global action.The data for years 2020, 2021, 2022, were removed from fitting to suppress the pandemic effects
The mean displacement as a function of time for each cluster was computed and is presented in Figure 9. This analysis clearly demonstrates that the clustering procedure effectively captures information on how country communities are progressing toward the ideal point. In the graph, the continuous lines represent the best fit to a second-order equation in time.
Table C.1 provides the actual coefficients for the clusters identified in Table B.1. Although any extrapolation based on these equations carries significant risk, it is possible to estimate the year by which, on average, the clusters will reach the ideal point. These estimations are included in the fifth column of Table C.1. The results indicate that the trajectories for all clusters extend well beyond 2030. The first region projected to approach the goals is Cluster 0, with an estimated year of 2048. According to current trends, Western developed countries are expected to meet the goals around the year 2085.
@@ 4 Discussion
In this study, we have shown that globally the SDGs show strong correlations. In most cases, these correlations are positive, which seems beneficial. In reference [21], where PCA was also used based on global data from 2000-2016, the authors found that synergies prevail among the SDGs, with SDG 10 (reducing inequality) being an exception. According to the results presented here, for the time span from 2000 to 2022, SDG 12 (responsible consumption and production) and SDG 13 (climate action) are strongly positively correlated with each other, but strongly negatively correlated with all other SDGs. This finding raises a potential red flag, indicating that the current global economic system may be working against the climate action goals. This inference is based on the correlation matrix derived from the goals of each country, suggesting that the vast network of global consumption and production is a contributing factor.
@@
In reference [22], the PCA is also used, but only, for the 35 countries of the OECD. The authors found that some countries perform significantly better compared to others, reflecting the heterogeneity within the OECD group, which is based on political agreements. The advantage of our study lies in the combined use of PCA and t-SNE, which allows for clustering based on real data differences, where each group of countries demonstrates similar behavior, primarily due to geographical, cultural, and socio-economic differences.
@@
In Figure 9, the trajectories of the mean distance to an ideal point for each cluster from Table B.1 are shown. All clusters are pursuing an ideal point, starting from different initial conditions. It is important to note that in 2020, the progress slowed down, most likely due to the COVID-19 pandemic.
There seems to be a relation among the standardized scores patterns and the dynamics. There are different rates of change. It appears that the wealthiest countries, in general, are the slowest to achieve the overall goal, while the less wealthy countries make faster progress. In any case, extrapolating the current tendencies (see Appendix C), no country would be reaching
the goals by 2030.
The patterns found in Fig. 6 could be interpreted as an imprint from some common policies implemented in the members of a cluster, that produce close results. It does not seem to be a coincidence that clusters reveal those patterns related to different trajectories in Fig. 9 that show
how a set of countries progresses in reaching the goals. Although the trajectories shown in the plot suggest that some clusters of countries may be on a path toward achieving the Sustainable Development Goals (SDGs) a few decades after the 2030 target, this optimistic outlook is likely to be disrupted by several significant global challenges. The main ones are the 1.5°C and 2°C global warming thresholds expected around 2030 and 2050, respectively [23]. Crossing these thresholds could exacerbate climate-related disasters, undermine ecosystems, and severely impact agricultural and water resources, making the achievement of the SDGs much more difficult. Additionally, the projected population growth to 9 billion by 2037 and 10 billion by 2055 will place increasing pressure on natural resources [24], which would intensify the competition and lead potentially to the depletion of critical resources such as freshwater and minerals by mid-century. The biodiversity crisis projected around 2050, if the current trends continue, could further delay progress by destabilizing the ecosystems which many human activities depend on [25]. Economic shifts, technological advances, and the persistent effects of the COVID-19 pandemic [26] also introduce uncertainties that could alter the course of sustainable development. Therefore, while the plotted extrapolated trajectories indicate a much delayed potential progress, the convergence toward the SDGs even decades later could be challenged by these obstacles, casting doubt on the feasibility of achieving these goals even then.
@@ Acknowledgments
Appendices
Appendix A Definition of SDG
@@
Fig A.1: Meaning of the 17 SDG according to the UN
Appendix B Countries in the different clusters
We show explicitly the names of the countries belonging to each cluster in Table B.1. Notice that cluster Number -1 is not really a cluster, it groups dots classified by DBSCAN as not belonging to any cluster since they lie in low density zone. Nevertheless, we include them for completeness. Also notice that few countries are in more that one cluster. This means that during the time span of the analysis they changed their situation
Table B.1: Countries grouped into clusters based on their SDG trajectories. The data underwent a sequential analysis: first, Principal Component Analysis (PCA) was utilized for dimensionality reduction. Then, t-Distributed Stochastic Neighbor Embedding (t-SNE) was applied to further map the data into a two-dimensional space, Finally, the DBSCAN clustering algorithm was used to group countries based on the resulting patterns.
@@ Appendix B Countries in the different clusters
Appendix C Dynamics
In here we show the values of the coefficients of the second order equation used to fit the dynamical behavior of the goals. According to this the prediction of the dates to achieve the ideal score is also shown in Table C.1.
@@ Appendix C Dynamics
Table C.1: Best fitting coefficients of equation $r(t)=a+bt+ct^{2}$ for each cluster of countries obtained in Table B.1. The last column is the extrapolated time to plenty achieve all the goals. Fitting was done without 2020, 2021 and 2022 data, to suppress the effect of COVID pandemic.
Appendix D The t-SNE dimensional reduction method.
@@
D.0.1 Probability Distributions in High Dimensions
@@
D.0.2 Mapping to Lower Dimensions

### v2

@@ 3.5 Cluster dynamics
The evolution of the SDGs of each country can be tracked over time using the full 17-dimensional vector of indicators. Each component of this vector has an indicator valued between 0 and 1, representing the degree to which the country has achieved a specific goal. This makes the vector point formed by these 17 values, denoted as $\vec{r}_{i}$, a unique vector that represents the ideal state that is being pursued. Consequently, for any given country at any point in time, the Euclidean distance to $\vec{r}_{i}$ can be monitored and visualized. Furthermore, by using the clusters described in S1 Table 2.1 (in Supplementary material S1 section), it is possible to calculate the distribution of distances to the ideal point for each year. In Figure 9 we show the probability density functions for the six clusters in the years 2000, 2010, and 2020. This figure reveals a continuous shift in the distribution functions which maintain approximately a Gaussian shape.
The Gaussian like distribution for the distances to the ideal goals within each cluster as seen in Figure  9 suggests that the countries grouped together share a similar pattern in their progress towards the SDGs. Moreover, the fact that these clusters are Gaussian indicates that within each of them, the countries are relatively homogeneous in terms of their distance to the ideal SDG foals. This homogeneity implies that the factors influence their progress are consistent for the countries in each cluster and that naturally group together based on shared challenges or opportunities to reach the SDGs. These clusters likely reflect countries that, despite their geographical or cultural differences, face similar systemic issues or have adopted similar strategies in their progress towards the SDGs.
@@
Fitted probability density functions for the distribution of distances to the ideal point for three different years and for each cluster on S1 Table 3.1 (in Supplementary material S1 section). The colors for the clusters are the same as in Figure 6. The Gaussian like distributions observed indicate a central tendency and homogeneity within each cluster with respect to their progress toward the SDGs.
@@
(A) Trajectories of clusters obtained in S1 Table 3.1 (in Supplementary material S1 section) toward the ideal achievement of the Sustainable Development Goals (SDGs) over time. Each line represents the projected average distance to the ideal point for a cluster of countries, with actual data points marked by dots, tracking progress from 2000 to 2022. (B) Extrapolation of same lines up to year 2100. The vertical dotted line at 2030 marks the point by which the countries of each cluster were expected to meet their SDG targets.
The mean displacement as a function of time for each cluster was calculated and its trajectory was fitted to the second order equation of time, as depicted in Figure 10. This analysis demonstrates that the clustering procedure effectively captures information on how the clusters of countries progress towards the ideal target point at 2030. S1 Table 2.1 (in Supplementary material S1 section) provides the actual coefficients for the clusters identified in S1 Table 3.1 (in Supplementary material S1 section). Although any extrapolation based on these equations carries significant risk, it is possible to estimate the year by which, on average, the clusters will reach the ideal point. These estimates are included in the fifth column of S1 Table 3.1 (in Supplementary material S1 section). The results of extrapolation in Figure 10B indicate that the trajectories of all clusters extend well beyond the 2030 target point.
@@ 4 Discussion
In this study, we have shown that globally the SDGs show strong correlations. In most cases, these correlations are positive, which is considered beneficial. In reference [23], where the PCA was also used based on global data from 2000-2016, the authors found that synergies prevail among the SDGs, with SDG 10 (reducing inequality) being an exception. According to the results presented here, for the time span from 2000 to 2022, SDG 12 (responsible consumption and production) and SDG 13 (climate action) are strongly positively correlated with each other, but strongly negatively correlated with all the other SDGs. This finding raises a potential red flag, indicating that the current global economic system may be working against the climate action goals. This inference is based on the correlation matrix derived from the goals of each country, suggesting that the vast network of global consumption and production is a contributing factor.
There is a general agreement with reference [8] where they found that SDG 1 (No poverty) has synergetic relationship with most of the other goals, while SDG 12 is the goal commonly associated with trade-offs. Reference [11] suggest that further research is needed to clarify the systemic roles of some SDGs (e.g., SDGs 12) to understand causal relationships. However, it is important to mention that it is difficult to compare exactly with previous global results, because the number of countries, the time span considered and the geographical regions [24] vary.
@@
In reference [25], the PCA is also used, but only, for 35 countries of the OECD. The authors found that some countries perform significantly better compared to others, reflecting the heterogeneity within the OECD group, which is based on political agreements. The advantage of our study lies in the combined use of PCA and t-SNE, which leads to an original clustering, based on real data. It is important to notice that clusters were obtained using solely SDG scores, showing that groups of countries having akin SDG pattens exhibit similar behavior, primarily due to geographical, cultural, and socio-economic features.
Based on our results it is possible to provide some general insights. African and Asian countires (Clusters 1 and 4), which are the ones having the lowest GDP/c, \$1,464.44 and \$3,941.89, respectively, show low SDG scores for Goals 1 to 11, 16 and 17 but posses the best performance with respect to responsible consumption and production and climate action (Goals 12 and 13); in contrast, the developed countries (Clusters 3 and 5), having the highest GDP/c, \$44,846.74 and \$43,524.00, respectively, show good achievements for Goals 1 to 11, 16 and 17 but have problems to improve their scores for Goals 12 and 13. The positive correlation between Goals 12 and 13 is very clear here, responsible consumption and production helps prevent climate change, but some other the economic factors seem to act against these goals. Latin America (Cluster 2), having the GDP/c \$10,328.50, has a different SDG pattern, showing the SDG scores above the mean for Goals 1 to 9, 11 to 14 and 17, and the lowest score for Goal 10, i.e., related to actions to reduce inequalities.
@@
In Figure 10, the trajectories of the mean distance to an ideal point for each cluster from S1 Table 3.1 (in Supplementary material S1 section) are shown. All clusters are pursuing an ideal point, starting from different initial conditions. It is important to note that in 2020, the progress slowed down, most likely due to the COVID-19 pandemic.
There seems to be a relation among the standardized scores patterns and the dynamics. There are different rates of change. It appears that the wealthiest countries, in general, are the slowest to achieve the overall goal, while the less wealthy countries make faster progress. In any case, extrapolating the current tendencies (see Appendix C), no country would be reaching the goals by 2030.
The patterns found in Fig. 7 could be interpreted as an imprint from some common policies implemented in the members of a cluster, that produce close results. It does not seem to be a coincidence that clusters reveal those patterns related to different trajectories in Fig. 10 that show how a set of countries progresses in reaching the goals. Although the trajectories shown in the plot suggest that some clusters of countries may be on a path toward achieving the Sustainable Development Goals (SDGs) a few decades after the 2030 target, this optimistic outlook is likely to be disrupted by several significant global challenges. The main ones are the 1.5°C and 2°C global warming thresholds expected around 2030 and 2050, respectively [26]. Crossing these thresholds could exacerbate climate-related disasters, undermine ecosystems, and severely impact agricultural and water resources, making the achievement of the SDGs much more difficult.
Additionally, the projected population growth to 9 billion by 2037 and 10 billion by 2055 will place increasing pressure on natural resources [27], which would intensify the competition and lead potentially to the depletion of critical resources such as freshwater and minerals by mid-century. The biodiversity crisis projected around 2050, if the current trends continue, could further delay progress by destabilizing the ecosystems which many human activities depend on [28]. Economic shifts, technological advances, and the persistent effects of the COVID-19 pandemic [29] also introduce uncertainties that could alter the course of sustainable development. Therefore, while the plotted extrapolated trajectories indicate a much delayed potential progress, the convergence toward the SDGs even decades later could be challenged by these obstacles, casting doubt on the feasibility of achieving these goals even then.
@@ Acknowledgments
Supplementary material
6 Definition of SDG
@@
Fig 6.1: Meaning of the 17 SDG according to the UN
7 Countries in the different clusters
We show explicitly the names of the countries belonging to each cluster in S1 Table 7.1. Notice that cluster Number -1 is not really a cluster, it groups dots classified by DBSCAN as not belonging to any cluster since they lie in low density zone. Nevertheless, we include them for completeness. Also notice that few countries are in more that one cluster. This means that during the time span of the analysis they changed their situation
Table 7.1: Countries grouped into clusters based on their SDG trajectories. The data underwent a sequential analysis: first, Principal Component Analysis (PCA) was utilized for dimensionality reduction. Then, t-Distributed Stochastic Neighbor Embedding (t-SNE) was applied to further map the data into a two-dimensional space, Finally, the DBSCAN clustering algorithm was used to group countries based on the resulting patterns.
@@ 7 Countries in the different clusters
8 Dynamics
In here we show the values of the coefficients of the second order equation used to fit the dynamical behavior of the goals. According to this the prediction of the dates to achieve the ideal score is also shown in S1 Table 8.1.
@@ 8 Dynamics
Table 8.1: Best fitting coefficients of equation $r(t)=a+bt+ct^{2}$ for each cluster of countries obtained in S1 Table 7.1. The last column is the extrapolated time to plenty achieve all the goals. Fitting was done without 2020, 2021 and 2022 data, to suppress the effect of COVID pandemic.
9 The t-SNE dimensional reduction method.
@@
9.1 Probability Distributions in High Dimensions
@@ 9.1.1 Perplexity
9.2 Mapping to Lower Dimensions